OpenStudy (anonymous):

In set builder notation how do you write the solutions of 5+ d < 8 -

OpenStudy (anonymous):

\[5+ d \le 8\]

OpenStudy (anonymous):

.

OpenStudy (anonymous):

Well, \(5+d \leq 8\). So \(d\leq3\) So d can be anything within the space 3 and below to negative infinity so in set notation it is: \[\{x\in R\phantom{.}|\phantom{.}-\infty\lt d \leq 3\}\] or \[\{x\in R\phantom{.}|\phantom{.}d\in (-\infty,3\phantom{.}\big]\phantom{.}\}\]